# Compact stellar systems hosting an intermediate mass black hole: magnetohydrodynamic study of inflow-outflow dynamics

### v1

@@ 3.1 Radial profiles
To more quantitatively understand accretion dynamics, we present radial profiles of fluid quantities averaged over angles and time. To do this, we first interpolate the data onto a spherical grid ($r,\theta,\varphi$) logarithmically spaced in radius $r$, where the polar $\theta$-axis is aligned with the $z$-axis of the simulation. Then we define the time (from $t_{\mathrm{min}}$ to $t_{\mathrm{max}}$) and angle average of a quantity $A$ as
@@
Figures 3c and 3d present the radial profiles of radial velocity, $v_{\rm r}$, and temperature, $T$, respectively, demonstrating that these quantities are largely unaffected by the stellar wind configuration (at least for our assumption of circular orbits). In the inner region ($r\lesssim 10^{-3}$ pc), the temperature radial profile is similar to that of the density, $T\propto r^{-1}$, as expected for an approximately virial temperature. The temperature gradually decreases to $\sim 10^{6}$ K at the outer domain boundary. A minor outlier from these trends is the magnetized stellar wind simulation with $\beta_{\rm w}=100$. Magnetic fields partially slow the outflow, and as a result, the winds reach slightly lower radial velocities. Dissipation of magnetic fields also contributes additional heating, leading to marginally higher temperatures.
@@
Fig. 5 plots radial profiles of the time and angle-averaged specific angular momentum, $\mathbf{l}=\mathbf{r}\times\mathbf{v}$, where $\mathbf{r}$ and $\mathbf{v}$ are the position and velocity vectors, respectively, for both the magnitude of the average and the average of the magnitude scaled to the Keplerian specific angular momentum, $l_{\mathrm{kep}}=\sqrt{GM_{\bullet}r}$. These curves show that individual gas elements injected into the domain by the stellar winds generally have a super-Keplerian amount of angular momentum, meaning that they tend to move outwards in radius. The net angular momentum, on the other hand, is quite low in comparison because there is a significant amount of cancellation due to the lack of a consistent direction for the angular momentum. This is true even for the Model III disk-like stellar configuration, primarily because the wind speeds are higher than the orbital speeds. For distances from the black hole within the feeding region, fluid elements tend to have sub-Keplerian angular momentum ($\langle|\mathbf{l}|\rangle\sim 0.3$–0.4 $l_{\rm kep}$ for the hydrodynamic simulations), implying that only the lowest angular momentum gas from the stellar winds falls inwards. Since the wind speeds are relatively large, this gas corresponds with either low impact parameter trajectories (i.e., gas that is blown almost directly at the black hole) or with gas that loses its angular momentum through wind-wind shocks. Again, in this region, the magnitude net angular momentum is much lower than the average magnitude $|\langle\mathbf{l}\rangle|\lesssim 0.07$ $l_{\rm kep}$, implying the lack of coherent accretion structure. The simulation with magnetic fields has the lowest $\langle|\mathbf{l}|\rangle$, $\lesssim 0.2l_{\rm kep}$, because magnetic forces tend to slow the gas.
@@
Finally, for the magnetized wind simulation, Fig. 5 presents angle and time-averaged radial profiles of plasma beta ($\beta=2P/B^{2}$) for different averaging methods: $\langle\beta\rangle$ (solid line), $\langle P\rangle_{\rho}/\langle B^{2}/2\rangle_{\rho}$ (dashed line), and $\langle\tilde{\beta}\rangle_{\rho}=\langle P+\rho v^{2}\rangle_{\rho}/\langle B^{2}/2\rangle_{\rho}$ (dotted line), where we include the gas ram pressure along with the thermal pressure. Of these three curves, $\langle\tilde{\beta}\rangle_{\rho}$ is best representative of the relative strength of magnetic energy/pressure relative to purely hydrodynamic energy/pressure. We find that $\langle\tilde{\beta}\rangle_{\rho}\sim 10$ for essentially all radii, implying that hydrodynamic forces dominate over magnetic ones. This is consistent with the radial profiles shown in Fig. 3, where the MHD simulation shows only moderate differences with the hydrodynamic simulations. $\langle P\rangle_{\rho}/\langle B^{2}/2\rangle_{\rho}$ is generally lower than $\langle\tilde{\beta}\rangle_{\rho}$ because the ram pressure is always significant when compared with the thermal pressure. In the region containing the stellar winds ($r\sim 10^{-3}\mathrm{pc}$$10^{-2}\mathrm{pc}$), $\langle P\rangle_{\rho}/\langle B^{2}/2\rangle_{\rho}$ is particularly low ($\sim$1–3) compared to $\langle\tilde{\beta}\rangle_{\rho}$ because most of the hydrodynmical energy is contained in the kinetic energy of the winds, which have low thermal pressure.
@@ 4.1 X-ray Surface brightness density
To evaluate the detectability of the IMBH embedded within the cluster, we calculate the X-ray surface brightness density, $S_{\mathrm{X}}$. We use Model II spherical as our fiducial model, since the wind chemical composition here best aligns with the realistic expectation, and we see from Fig. 2-5 that all models are qualitatively similar. First, we use spex as described in Sect. 2.1 to calculate the cooling function, $\Lambda$, except we consider only the contributions from frequencies corresponding to the Chandra merged hard-band ($2-7$ keV) (Evans et al., 2010), denoting this as $\Lambda_{\mathrm{X}}$. Then we transform our simulation data to cylindrical coordinates ($s,z,\varphi$, with $z$ being the line of sight) and integrate along the line of sight to obtain
@@
where $\eta$ is the radiative efficiency (assumed to be 0.1) and $\sigma_{\rm T}$ is the Thomson cross-section. We highlight specific spectral bands and include an approximate Sgr A* SED calculated using the same model for a “typical” accretion rate of $2\times 10^{-8}\,M_{\odot}\,\mathrm{yr}^{-1}$, corresponding to an Eddington ratio of $\lambda_{\mathrm{Edd}}=2.25\times 10^{-7}$ (Marrone et al., 2007). Compared to Sgr A*, the IMBH’s radiative flux is roughly three orders of magnitude lower across the spectrum, with its peak slightly shifted from the far-infrared (FIR) to the mid-infrared (MIR).
@@
where $r_{\mathrm{in}}$ is the radius of the inner boundary. Fig. 8 illustrates the time evolution of the X-ray luminosity for the ADAF SED model associated with the IMBH and the large-scale emission from the IMBH, computed from Eq. (8) integrated over the the central $10^{-4}$ pc encompassing the black hole (left), the large-scale flow modeled by our simulations, computed from Eq. (8) integrated over the whole domain (middle), and all lines combined on the same plot (right). The luminosity produced by the large-scale flow is approximately two orders of magnitude greater than both the ADAF luminosity and the large-scale emission associated with the IMBH, reaffirming that stellar wind collisions dominate the X-ray spectrum.
@@ 4.2 X-ray data analysis
To compare the simulated X-ray signal with actual X-ray data of the IRS 13E source, we use archival observations obtained by the Chandra X-ray Observatory. In total, we analyze 128 non-grating ACIS-S and ACIS-I Chandra Observational IDs (OBSIDs) with Sgr A* in the aim-point spanning almost 25 years. All observations are reprocessed using the chandra_repro script (ciao 4.17; Fruscione et al., 2006) and the latest calibration files (caldb 4.11.6).
@@
To obtain the best possible match between observations, we perform relative astrometric corrections using the fine_astro script. As a reference, we use the OBSID 11843, which provides long enough exposure (78 ks) with a minimum of bright transient sources close to the chip aim-point, which would complicate source-matching with other OBSIDs. Before merging all OBSIDs, we further omit observations containing very bright transient sources in the vicinity of IRS 13E that would contaminate its signal, and we end up with 112 observations with a total cleaned exposure time of $\approx 3.6$ Ms ($\approx 2$ Ms for ACIS-S and $\approx 1.6$ for ACIS-I).
@@
When reprocessing the individual OBSIDs using the chandra_repro script, we use the default value for the pix_adj parameter (pix_adj=EDSER), which allows us to leverage the sub-pixel Chandra resolution. We then produce a merged hard-band ($2-7$ keV) image binned to a pixel size of $\approx 0.125$ arcsec. This allows us to study the morphology of the X-ray source corresponding to IRS 13E.
@@
The simulated mock Chandra image is obtained by scaling the X-ray surface brightness map (Fig. 6) to a distance of $8$ kpc and then using the SOXS package (ZuHone et al., 2023) to simulate a mock Chandra ACIS-S [^1] Effective area corresponding to Cycle 22 was assumed. observation with 3.6 Ms exposure time. Based on the simulated event file, we then produce a hard-band image binned to match the pixel size of the real Chandra sub-pixel image ($\approx 0.125$ arcsec). The sub-pixel Chandra image is compared to the simulated image at $t=50$ yr in Fig. 9. We chose this time arbitrarily, since the simulated image does not change much throughout the simulation, with only the emission centroid moving slightly depending on where the winds happen to be colliding the most. Given the uncertainties in the stellar winds parameters, the qualitative agreement between the Chandra image and the mock image from our simulation is quite good. The simulation produces a slightly more spherical distribution of X-ray luminosity, which is not surprising because we assumed that all the stars were on spherical orbits. The Chandra image also has more extended, diffuse emission outside the brighter core than what we find. This could also be a result of our assumption of circular orbits; more eccentric orbits would occasionally bring the stellar winds out to larger radii and populate those regions with hot gas. Alternatively, our simulations could be fainter there because we do not include any stars or gas other than the six listed in Table 3.
@@ 4.3 Accretion rate and X-ray Variability
In order to isolate candidate periodicities in the accretion rate, we compute Lomb-Scargle periodograms (Lomb, 1976; Scargle, 1982), as detailed in Appendix B. Our simulations typically span 2340 years and we sample every 5 years, meaning that we can identify periodic signals in the time range from 10 to 1000 years. For the accretion rate in spherically distributed stellar systems, we detect a prominent periodicity peak at $\sim 110$ years for both low and high metallicity. In addition, peaks at $\sim 200$ years, $\sim 55$ years, and $\sim 70$ years are also pronounced, as shown in Figs. 13-15 (left panels). For the disk-like stellar configuration, the most prominent peak is at $\sim 186$ years, followed by peaks at $\sim 217$ and $\sim 100$ years, respectively, as shown in Fig. 16 (left panel).
These detected periodicities are induced by close encounters between stars, during which their strong stellar winds are compressed and shocked. Most of the shocked, denser gas flows out of the system while a small fraction gets accreted. In Appendix A, there is a comprehensive overview of the approaches of the stellar pairs, including minimum, mean, and maximum distances (in AU) and the associated recurrence timescales (in years). The most pronounced periods are associated with the star E4, which has the most compact orbit (at a distance of $4.06$ mpc). For the adopted spherical setup of stars, the closest approaches between stars E4 and E5.1 take place every 108 years, which is consistent with the most prominent periodicity peak in the inflow rate, compare Fig. 10 (left column) and Figs. 13-15. On the other hand, for the disk-like setup, the approaches between the same stars (E4 and E5.1) take place every 198 years, which is again consistent with the broader peak around 186-217 years in the periodogram (Fig. 16; compare with Fig. 10, right column). The closest approaches between star E4 and other stars take place with comparable periods of $\sim 100-200$ years (see Tables 4 and 5), which further affects the mass inflow rate; specific inflow peaks are strengthened or prolonged when the stellar approaches take place almost simultaneously. In addition, occasional rare approaches on the scale of $\sim 100-200\,{\rm AU}$ among other stars can further significantly modulate the inflow rate at some epochs. Altogether, this results in the presence of multiple periodicity peaks in the periodograms for all the runs.
@@
Finally, we compute Lomb-Scargle periodograms (Fig. 11) and PSDs (Fig. 12) for the X-ray light curves presented in §4.2. Because both the magnitude and the spectral shape of the X-ray portion of the ADAF model’s SED (Fig. 7) depend on the variable accretion rate, the time-variability is not one-to-one with the accretion rate. The Lomb-Scargle periodograms of the ADAF models display significant peaks at $\sim 210$ years and $\sim 54$ years, while the peak around $\sim 110$ years appears to be suppressed compared to the periodogram of the accretion rate in Fig. 13. On the other hand, the total X-ray luminosity, which is dominated by the shocked stellar winds, has a significant periodic component at $\sim 104$ years, corresponding to the periods of close approaches among the stellar pairs E4-E1 and E4-E5.1. In terms of power, the PSD for the ADAF model is significantly flatter than the PSD for the whole cluster (with power-law slopes of $\sim-0.53$ vs. $\sim-1.41$). This is consistent with the fact that there is more short-term variability on timescales of $<100$ years in the region closer to the inner boundary of the simulation than on larger scales of the cluster.
@@
where $\sigma^{2}$ is the variance of the luminosity, $\Delta^{2}$ is the mean square uncertainty associated with individual luminosities $L_{i}$, and $\overline{L}$ is the mean luminosity. Assuming one percent relative uncertainty for both the IMBH ADAF and the total (cluster) luminosities (see Fig. 8 left and middle panels), we obtain $F_{\rm var}\simeq 213\%$ for the IMBH’s ADAF X-ray luminosity, while we infer $F_{\rm var}\simeq 20\%$ for the X-ray luminosity of the whole cluster (for a luminosity uncertainty of $10\%$, $F_{\rm var}\simeq 212\%$ and $F_{\rm var}\simeq 17\%$ for the IMBH’s ADAF and the total luminosities, respectively). Hence, the cluster X-ray emission, whose variability is rather mild, masks a much more vigorous X-ray emission of the IMBH ADAF. This is also consistent with the flatter power spectral density of the ADAF variability ($\propto f^{-0.5}$) in comparison with PSD of the total luminosity ($\propto f^{-1.4}$), see Fig. 12. The quasiperiodic X-ray spikes associated with the IMBH with the typical timescale of $\sim 100$ years are reminiscent of order-of-magnitude quasiperiodic X-ray eruptions (QPEs) detected for several both active and quiescent galactic nuclei (Miniutti et al., 2019; Arcodia et al., 2021; Suková et al., 2024). However, in contrast, QPEs take place on the timescales five orders of magnitude smaller and the QPE X-ray properties are consistent with being associated with the soft thermal X-ray emission. Still, in future work, it is worth considering a compact association of wind-blowing stars orbiting around the massive black hole in a galactic nucleus as a potential model to address some of the QPE sources. Such a simulation would have to be a scaled-down version of the MHD simulations presented in this work with the characteristic stellar orbit radius of $r_{\rm QPE}=r_{\rm IRS13}(P_{\rm QPE}/P_{\rm IRS13})^{2/3}\sim 2\times 10^{-6}\,(r_{\rm IRS13}/4\times 10^{-3}\,{\rm pc})(P_{\rm QPE}/10\,\text{hours})^{2/3}(P_{\rm IRS13}/100\,\text{years})^{-2/3}{\rm pc}$ for $M=3\times 10^{4}\,M_{\odot}$ (IMBH), which corresponds to $r_{\rm QPE}\sim 5.4\times 10^{-6}\,{\rm pc}\sim 111(P_{\rm QPE}/10\,\text{hours})^{2/3}(M/10^{6}\,M_{\odot})^{-2/3}$ gravitational radii for the $M=10^{6}\,M_{\odot}$ supermassive black hole.
Figure 11: Lomb-Scargle (LS) periodograms of the X-ray light curves computed from the simulations. Left: The LS periodogram of the X-ray light curve produced by the near-horizon flow around the IMBH as calculated from an ADAF model. The highest peak is at $\sim 210$ years, while other prominent peaks are at $\sim 54$ and $\sim 83$ years. Dotted horizontal lines mark inferred confidence levels based on 10000 bootstrap realizations. Right: The LS periodogram of the total X-ray luminosity exhibits the highest peak at $\sim 104$ years. Other prominent peaks are at $\sim 202$, $\sim 271$, $\sim 55$, and $\sim 74$ years. Dotted horizontal lines stand for the confidence levels as in the left panel.

### v2

@@ 3.2 Radial profiles
We present radial profiles of fluid quantities averaged over angles and time to better understand accretion dynamics. To do this, we first interpolate the data onto a spherical grid ($r,\theta,\varphi$) logarithmically spaced in radius $r$, where the polar $\theta$-axis is aligned with the $z$-axis of the simulation. Then we define the time (from $t_{\mathrm{min}}$ to $t_{\mathrm{max}}$) and angle average of a quantity $A$ as
@@
Figures 3c and 3d present the radial profiles of radial velocity, $v_{\rm r}$, and temperature, $T$, respectively, demonstrating that these quantities are largely unaffected by the stellar wind configuration (at least for our assumption of circular orbits). In the inner region ($r\lesssim 10^{-3}$ pc), the temperature radial profile is similar to that of the density, $T\propto r^{-1}$, as expected for an approximately virial temperature. The temperature gradually decreases to $\sim 10^{6}$ K at the outer domain boundary. The magnetized stellar wind simulation with $\beta_{\rm w}=100$ is a minor outlier from these trends. Magnetic fields partially slow the outflow, so the winds reach slightly lower radial velocities. Dissipation of magnetic fields also contributes additional heating, leading to marginally higher temperatures.
@@ 3.3 Angular momentum and magnetic effects
Fig. 5 plots radial profiles of the time and angle-averaged specific angular momentum, $\mathbf{l}=\mathbf{r}\times\mathbf{v}$, where $\mathbf{r}$ and $\mathbf{v}$ are the position and velocity vectors, respectively, for both the magnitude of the average and the average of the magnitude scaled to the Keplerian specific angular momentum, $l_{\mathrm{kep}}=\sqrt{GM_{\bullet}r}$. These curves show that individual gas elements injected into the domain by the stellar winds generally have a super-Keplerian amount of angular momentum, meaning they tend to move outwards in radius. On the other hand, the net angular momentum is relatively low in comparison because there is a significant amount of cancellation due to the lack of a consistent direction for the angular momentum. This is true even for the Model III disk-like stellar configuration, primarily because the wind speeds are higher than the orbital speeds. For distances from the black hole within the feeding region, fluid elements tend to have sub-Keplerian angular momentum ($\langle|\mathbf{l}|\rangle\sim 0.3$–0.4 $l_{\rm kep}$ for the hydrodynamic simulations), implying that only the lowest angular momentum gas from the stellar winds falls inwards. Since the wind speeds are relatively large, this gas corresponds with either low impact parameter trajectories (i.e., gas blown almost directly at the black hole) or gas that loses its angular momentum through wind-wind shocks. Again, in this region, the magnitude of the net angular momentum is much lower than the average magnitude $|\langle\mathbf{l}\rangle|\lesssim 0.07$ $l_{\rm kep}$, implying the lack of coherent accretion structure. The simulation with magnetic fields has the lowest $\langle|\mathbf{l}|\rangle$, $\lesssim 0.2l_{\rm kep}$, because magnetic forces tend to slow the gas.
@@
Finally, for the magnetized wind simulation, Fig. 5 presents angle and time-averaged radial profiles of plasma beta ($\beta=2P/B^{2}$) for different averaging methods: $\langle\beta\rangle$ (solid line), $\langle P\rangle_{\rho}/\langle B^{2}/2\rangle_{\rho}$ (dashed line), and $\langle\tilde{\beta}\rangle_{\rho}=\langle P+\rho v^{2}\rangle_{\rho}/\langle B^{2}/2\rangle_{\rho}$ (dotted line), where we include the gas ram pressure along with the thermal pressure. Of these three curves, $\langle\tilde{\beta}\rangle_{\rho}$ best represents the relative strength of magnetic energy/pressure relative to purely hydrodynamic energy/pressure. We find that $\langle\tilde{\beta}\rangle_{\rho}\sim 10$ for all radii implies that hydrodynamic forces dominate over magnetic ones. This is consistent with the radial profiles shown in Fig. 3, where the MHD simulation shows only moderate differences with the hydrodynamic simulations. $\langle P\rangle_{\rho}/\langle B^{2}/2\rangle_{\rho}$ is generally lower than $\langle\tilde{\beta}\rangle_{\rho}$ because the ram pressure is always significant when compared with the thermal pressure. In the region containing the stellar winds ($r\sim 10^{-3}\mathrm{pc}$$10^{-2}\mathrm{pc}$), $\langle P\rangle_{\rho}/\langle B^{2}/2\rangle_{\rho}$ is particularly low ($\sim$1–3) compared to $\langle\tilde{\beta}\rangle_{\rho}$ because most of the hydrodynamical energy is contained in the kinetic energy of the winds, which have low thermal pressure.
@@ 4.1 X-ray Surface brightness density
To evaluate the detectability of the IMBH embedded within the cluster, we calculate the X-ray surface brightness density, $S_{\mathrm{X}}$. We use Model II spherical as our fiducial model, since the wind chemical composition here best aligns with the realistic expectation, and we see from Fig. 2-5 that all models are qualitatively similar. First, we use spex as described in Sect. 2.1 to calculate the cooling function, $\Lambda$, except we consider only the contributions from frequencies corresponding to the Chandra merged hard-band (272-7 keV) (Evans et al., 2010), denoting this as $\Lambda_{\mathrm{X}}$. Then we transform our simulation data to cylindrical coordinates ($s,z,\varphi$, with $z$ being the line of sight) and integrate along the line of sight to obtain
@@
where $\eta$ is the radiative efficiency (assumed to be 0.1) and $\sigma_{\rm T}$ is the Thomson cross-section. We highlight specific spectral bands and include an approximate Sgr A* SED calculated using the same model for a ”typical” accretion rate of $2\times 10^{-8}\,M_{\odot}\,\mathrm{yr}^{-1}$, corresponding to an Eddington ratio of $\lambda_{\mathrm{Edd}}=2.25\times 10^{-7}$ (Marrone et al., 2007). Compared to Sgr A*, the IMBH’s radiative flux is roughly three orders of magnitude lower across the spectrum, with its peak slightly shifted from the far-infrared (FIR) to the mid-infrared (MIR).
@@
where $r_{\mathrm{in}}$ is the radius of the inner boundary. Fig. 8 illustrates the time evolution of the X-ray luminosity for the ADAF SED model associated with the IMBH and the large-scale emission from the IMBH, computed from Eq. (8) integrated over the the central $10^{-4}$ pc encompassing the black hole (left), the large-scale flow modeled by our simulations, computed from Eq. (8) integrated over the whole domain (middle), and all lines combined on the same plot (right). The luminosity produced by the large-scale flow is approximately two orders of magnitude greater than the ADAF luminosity and the large-scale emission associated with the IMBH, reaffirming that stellar wind collisions dominate the X-ray spectrum.
@@ 4.2 X-ray data analysis
To compare the simulated X-ray signal with actual X-ray data of the IRS 13E source, we use archival observations obtained by the Chandra X-ray Observatory. We analyzed 128 non-grating ACIS-S and ACIS-I Chandra Observational IDs (OBSIDs) with Sgr A* in the aim-point spanning almost 25 years. All observations are reprocessed using the chandra_repro script (ciao 4.17; Fruscione et al., 2006) and the latest calibration files (caldb 4.11.6).
@@
We perform relative astrometric corrections using the fine_astro script to obtain the best possible match between observations. As a reference, we use the OBSID 11843, which provides long enough exposure (78 ks) with a minimum of bright transient sources close to the chip aim-point, which would complicate source-matching with other OBSIDs. Before merging all OBSIDs, we further omit observations containing very bright transient sources in the vicinity of IRS 13E that would contaminate its signal, and we end up with 112 observations with a total cleaned exposure time of $\approx 3.6$ Ms ($\approx 2$ Ms for ACIS-S and $\approx 1.6$ for ACIS-I).
@@
When reprocessing the individual OBSIDs using the chandra_repro script, we use the default value for the pix_adj parameter (pix_adj=EDSER), which allows us to leverage the sub-pixel Chandra resolution. We then produce a merged hard-band (272-7 keV) image binned to a pixel size of $\approx 0.125$ arcsec. This allows us to study the morphology of the X-ray source corresponding to IRS 13E.
@@
The simulated mock Chandra image obtained by scaling the X-ray surface brightness map (Fig. 6) to a distance of $8$ kpc and then using the SOXS package (ZuHone et al., 2023) to simulate a mock Chandra ACIS-S [^1] Effective area corresponding to Cycle 22 was assumed. observation with 3.6 Ms exposure time. Based on the simulated event file, we then produce a hard-band image binned to match the pixel size of the real Chandra sub-pixel image ($\approx 0.125$ arcsec). We compare the sub-pixel Chandra image to the simulated image at $t=50$ yr in Fig. 9. We chose this time arbitrarily, since the simulated image does not change much throughout the simulation, with only the emission centroid moving slightly depending on where the winds collide the most. Given the uncertainties in the stellar wind parameters, the qualitative agreement between the Chandra image and the mock image from our simulation is quite good except for minor differences. In Appendix C and in particular in Fig. 17, we compare the surface-brightness profiles between the observed image and the simulated one. The simulation produces a slightly more spherical distribution of X-ray luminosity, which we demonstrate by evaluating the asymmetry parameter ($A=0.13$ for the simulated image vs. $A=0.21$ for the observed one). This result is not surprising because we assumed that all the stars were on spherical orbits. More importantly, the Chandra image also contains an extended, diffuse emission outside the brighter core. In addition, there are additional background/foreground X-ray point sources in the IRS 13E field. Our simulations do not contain any more distant sources, and the simulated image is thus based entirely on the six stars listed in Table 3 that orbit the IMBH. However, apart from the differences in the tails of the X-ray surface-brightness profile, both profiles are qualitatively consistent with the uncertainties. They closely follow and are slightly wider than the Chandra point-spread function (PSF). From a quantitative point of view, the total flux/luminosity of IRS 13E and the simulated cluster differ significantly, with the simulated cluster being more luminous by an order of magnitude. The agreement could be reached by fine-tuning, e.g., the WR stars’ mass-loss rates and wind velocities within the uncertainties. Since we did not seek an exact model for IRS 13E at this point, we will explore the parameter space of the cluster in relation to IRS 13E in the upcoming study in more detail.
@@ 4.3 Accretion rate and X-ray Variability
In order to isolate candidate periodicities in the accretion rate, we compute Lomb-Scargle periodograms (Lomb, 1976; Scargle, 1982), as detailed in Appendix B. Our simulations typically span 2340 years, and we sample every 5 years, meaning that we can identify periodic signals in the time range from 10 to 1000 years. For the accretion rate in spherically distributed stellar systems, we detect a prominent periodicity peak at $\sim 110$ years for both low and high metallicity. In addition, peaks at $\sim 200$ years, $\sim 55$ years, and $\sim 70$ years are also pronounced, as shown in Figs. 13-15 (left panels). For the disk-like stellar configuration, the most prominent peak is at $\sim 186$ years, followed by peaks at $\sim 217$ and $\sim 100$ years, respectively, as shown in Fig. 16 (left panel).
These detected periodicities induced close encounters between stars, during which their strong stellar winds are compressed and shocked. The most shocked, denser gas flows out of the system while a small fraction accretes onto the black hole. In Appendix A, there is a comprehensive overview of the approaches of the stellar pairs, including minimum, mean, and maximum distances (in AU) and the associated recurrence timescales (in years). The most pronounced periods are associated with the star E4, which has the most compact orbit (at a distance of $4.06$ mpc). For the adopted spherical setup of stars, the closest approaches between stars E4 and E5.1 take place every 108 years, which is consistent with the most prominent periodicity peak in the inflow rate, compare Fig. 10 (left column) and Figs. 13-15. On the other hand, for the disk-like setup, the approaches between the same stars (E4 and E5.1) take place every 198 years, which is again consistent with the broader peak around 186-217 years in the periodogram (Fig. 16; compare with Fig. 10, right column). The closest approaches between star E4 and other stars take place with comparable periods of $\sim 100-200$ years (see Tables 4 and 5), which further affects the mass inflow rate; specific inflow peaks are strengthened or prolonged when the stellar approaches take place almost simultaneously. In addition, occasional rare approaches on the scale of $\sim 100-200\,{\rm AU}$ among other stars can further significantly modulate the inflow rate at some epochs. Altogether, this results in multiple periodicity peaks for all the runs in the periodograms.
@@
Finally, we compute Lomb-Scargle periodograms (Fig. 11) and PSDs (Fig. 12) for the X-ray light curves presented in Sect. 4.2. Because the magnitude and the spectral shape of the X-ray portion of the ADAF model’s SED (Fig. 7) depend on the variable accretion rate, the time-variability is not one-to-one with the accretion rate. The Lomb-Scargle periodograms of the ADAF models display significant peaks at $\sim 210$ years and $\sim 54$ years. In contrast, the peak around $\sim 110$ years appears to be suppressed compared to the periodogram of the accretion rate in Fig. 13. On the other hand, the total X-ray luminosity, which dominated the shocked stellar winds, has a significant periodic component at $\sim 104$ years, corresponding to the periods of close approaches among the stellar pairs E4-E1 and E4-E5.1. In terms of power, the PSD for the ADAF model is significantly flatter than the PSD for the whole cluster (with power-law slopes of $\sim-0.53$ vs. $\sim-1.41$). This is consistent with the fact that there is more short-term variability on timescales of <100<100 years in the region closer to the inner boundary of the simulation than on larger scales of the cluster.
@@
where $\sigma^{2}$ is the variance of the luminosity, $\Delta^{2}$ is the mean square uncertainty associated with individual luminosities $L_{i}$, and $\overline{L}$ is the mean luminosity. Assuming one percent relative uncertainty for both the IMBH ADAF and the total (cluster) luminosities (see Fig. 8 left and middle panels), we obtain $F_{\rm var}\simeq 213\%$ for the IMBH’s ADAF X-ray luminosity, while we infer $F_{\rm var}\simeq 20\%$ for the X-ray luminosity of the whole cluster (for a luminosity uncertainty of $10\%$, $F_{\rm var}\simeq 212\%$ and $F_{\rm var}\simeq 17\%$ for the IMBH’s ADAF and the total luminosities, respectively). Hence, the cluster X-ray emission, whose variability is mild, masks a much more vigorous X-ray emission of the IMBH ADAF. This is also consistent with the flatter power spectral density of the ADAF variability ($\propto f^{-0.5}$) in comparison with the PSD of the total luminosity ($\propto f^{-1.4}$), see Fig. 12. The quasiperiodic X-ray spikes associated with the IMBH with the typical timescale of $\sim 100$ years are reminiscent of order-of-magnitude quasiperiodic X-ray eruptions (QPEs) detected for several both active and quiescent galactic nuclei (Miniutti et al., 2019; Arcodia et al., 2021; Suková et al., 2024). However, in contrast, QPEs take place on timescales five orders of magnitude smaller, and the QPE X-ray properties are consistent with being associated with the soft thermal X-ray emission. Still, in future work, it is worth considering a compact association of wind-blowing stars orbiting around the massive black hole in a galactic nucleus as a potential model to address some of the QPE sources. Such a simulation would have to be a scaled-down version of the MHD simulations presented in this work with the characteristic stellar orbit radius of $r_{\rm QPE}=r_{\rm IRS13}(P_{\rm QPE}/P_{\rm IRS13})^{2/3}\sim 2\times 10^{-6}\,(r_{\rm IRS13}/4\times 10^{-3}\,{\rm pc})(P_{\rm QPE}/10\,\text{hours})^{2/3}(P_{\rm IRS13}/100\,\text{years})^{-2/3}{\rm pc}$ for $M=3\times 10^{4}\,M_{\odot}$ (IMBH), which corresponds to $r_{\rm QPE}\sim 5.4\times 10^{-6}\,{\rm pc}\sim 111(P_{\rm QPE}/10\,\text{hours})^{2/3}(M/10^{6}\,M_{\odot})^{-2/3}$ gravitational radii for the $M=10^{6}\,M_{\odot}$ supermassive black hole.
Figure 11: Lomb-Scargle (LS) periodograms of the X-ray light curves computed from the simulations. Left: The LS periodogram of the X-ray light curve produced by the near-horizon flow around the IMBH as calculated from an ADAF model. The highest peak is at $\sim 210$ years, while other prominent peaks are at $\sim 54$ and $\sim 83$ years. Dotted horizontal lines mark inferred confidence levels based on 10000 bootstrap realizations. Right: The LS periodogram of the total X-ray luminosity exhibits the highest peak at $\sim 104$ years. Other prominent peaks are at $\sim 202$, $\sim 271$, $\sim 55$, and $\sim 74$ years. Dotted horizontal lines stand for the confidence levels, as in the left panel.